# Deep Learning-based Prediction of Clinical Trial Enrollment with Uncertainty Estimates

### v1

@@ 4.3.2 Deep Poisson-Gamma
Table 4: The performance of the Deep Poisson-Gamma model in comparison with the “Filtering-and-Fitting” approach in terms of trial duration prediction. The unit of trial duration is month.
Filtering-and-Fitting Deep Poisson-Gamma
MAE ($\downarrow$) 10.5510.5510.55 $\mathbf{7.52}$
MedAE ($\downarrow$) 8.708.708.70 $\mathbf{5.59}$
6-months coverage [%] ($\uparrow$) 14.9414.9414.94 $\mathbf{32.18}$
Inference time [s] ($\downarrow$) 8.748.748.74 $\mathbf{0.07}$
As it can be seen in Table 4, our Deep Poisson-Gamma model outperforms the filtering-and-fitting approach on all metrics. More precisely, the former achieves an MAE of 7.527.527.52 months, a three-month error reduction from the latter. In terms of 6-month coverage accuracy, the proposed model performs twice as well as the filtering-and-fitting approach. Moreover, it can be seen that by leveraging machine learning paradigm, we can significantly reduce inference time. This is due to the fact that for the filtering-and-fitting approach, new distributions need to be fit on similar studies for each trial. This process can be time-consuming as what can be seen as “training” is part of the inference process. Our approach leveraging machine learning, allows to separate the training and inference processes and thus greatly reduce the inference time.

### v2

@@ 4.3.2 Deep Poisson-Gamma
As it can be seen in Table LABEL:table:poisson_gamma_results, our Deep Poisson-Gamma model outperforms the filtering-and-fitting approach on all metrics. More precisely, the former achieves an MAE of $7.85$ months, a two-month error reduction from the latter. Moreover, it can be seen that by leveraging machine learning paradigm, we can significantly reduce inference time. This is due to the fact that for the filtering-and-fitting approach, new distributions need to be fit on similar studies for each trial. This process can be time-consuming as what can be seen as “training” is part of the inference process. Our approach leveraging machine learning, allows to separate the training and inference processes and thus greatly reduce the inference time.